# MediSwift: Efficient Sparse Pre-trained Biomedical Language Models

### v1

@@ Abstract
Vithursan Thangarasa† Mahmoud Salem Shreyas Saxena Kevin Leong Joel Hestness Sean Lie Cerebras Systems vithu@cerebras.net
@@ 1 Introduction
Despite these advancements and the clear benefits of domain-specific LLM pre-training, the computational demands and associated costs remain a significant challenge, highlighting the need for innovative solutions to make these powerful tools more accessible and efficient for a wide-range of applications. In response, a variety of techniques have been proposed aimed at mitigating the computational burden associated with training LLMs. These include sparse attention (Dao et al., 2022b; Jaszczur et al., 2021), quantized optimization (Tang et al., 2021), low-rank factorization (Lialin et al., 2023) and sequence-level curriculum learning (Li et al., 2022). Among these approaches, weight sparsity emerges as a particularly promising method, distinct from the others in its approach by setting a subset of model parameters to zero, thus reducing the FLOPs needed during training.
@@
However, the adoption of sparse training is limited by (1) the challenge of finding optimal sparsity patterns which retain the accuracy of the original dense model (Frankle and Carbin, 2018; Ma et al., 2022) and (2) its difficulties in accelerating on hardware optimized for dense computations (e.g., GPUs and TPUs) (Hooker, 2020). Additionally, sparse pre-training of LLMs typically leads to an accuracy loss from optimization challenges in sparse networks (Evci et al., 2019), but previous studies have shown that transitioning from sparse-to-dense training can effectively address these problems (Thangarasa et al., 2023a; Dao et al., 2022a). Our work aims to reduce training FLOPs by employing unstructured weight sparsity in domain-specific LLMs and recovering lost representational capacity by transitioning to dense weight matrices for fine-tuning on downstream tasks.
@@
Despite weight sparsity not being widely used in real-world applications, advancements in specialized software kernels (Gale et al., 2019; Elsen et al., 2019; Ashby et al., 2019; Wang, 2021; Tang et al., 2023) have facilitated inference acceleration with unstructured sparsity. Recent developments have shown that the benefits of unstructured weight sparsity can be fully harnessed on specialized hardware, such as the Cerebras CS-2 (Cerebras, 2023; Lie, 2022b), for LLM training. As sparse training techniques and hardware continue to co-evolve, we anticipate that the reductions in FLOPs will lead to realized sparse acceleration. The latest innovations in software and hardware (NeuralMagic, 2021; Gupta, 2024) are geared towards enabling the widespread adoption of unstructured weight sparsity, offering the potential to achieve higher compression ratios and practical speedups in terms of wall-clock time.
@@
Building on this momentum, we introduce MediSwift, a suite of biomedical language models (LLMs) available in three sizes: Med (302M), Large (510M), and XL (1.2B). These models are based on GPT-3 and pre-trained sparsely from scratch on biomedical texts, aimed at reducing the computational costs required for training. We explore the impact of applying 50% and 75% weight sparsity during pre-training, which results in a 2-2.5x reduction in the overall training FLOPs needed. Figure 1 summarizes the performance, where our dense and sparse MediSwift models noticeably outperform other language models up to 7B parameters. We demonstrate MediSwift’s capabilities through fine-tuning on established benchmarks for biomedical natural language processing (NLP) tasks (e.g., PubMedQA (Jin et al., 2019) for question answering and HoC (Baker et al., 2016) for document classification) showing significant improvements in the balance between efficiency and accuracy. Although previous research suggests that sparse pre-training may compromise model accuracy on downstream tasks, our approach incorporates dense fine-tuning with strategic soft prompting to effectively regain performance on specialized tasks. Specifically, MediSwift-XL (1.2B) sets a new state-of-the-art by reaching 76.8% accuracy, despite being 5.8x smaller than PMC-LlaMA. Moreover, MediSwift-XL models, pre-trained with 50% and 75% sparsity, surpass the performance of models of similar or greater sizes (e.g., MediSwift-XL at 75% sparsity outperforms BioMedLM while being almost 9x smaller). Our work not only highlights the potential for sparse pre-training to make LM training more economically viable but also sets a new benchmark for efficiency in domain-specific applications of LLMs. The key contributions are:
@@ 2.1 Autoregressive Language Modeling
Building upon our framework for dense pre-training, we introduce weight sparsity into the model, specifically to improve the computational efficiency. We achieve this by methodically reducing the number of active connections within each layer of the model by a predefined sparsity level , effectively rendering a portion of the network’s parameters inactive. This process yields a network with active parameters per layer, where denotes the original number of parameters in layer . The overall sparsity is quantified by , which represents the ratio of inactive parameters to the total parameter count of the initially dense model, calculated as . To apply sparsity effectively, we employ a binary mask to the model’s initial parameters , resulting in a sparse parameter set . This mask effectively segregates the parameters into active (1) and inactive (0) states, thereby establishing a sparsity-induced version of the language model that aims to minimize a slightly modified objective:
We leverage the GPT-3 (Brown et al., 2020a) architecture for the MediSwift biomedical language model, training it with the AdamW (Loshchilov and Hutter, 2017) optimizer on a carefully curated biomedical dataset following the objective shown in Eq. 1 for iterations to get parameters . This pre-trained model is then fine-tuned for specific supervised tasks in the biomedical domain.
@@ 2.2 Dense Fine-tuning and Soft Prompting
In this section, we detail the adaptation of our pre-trained MediSwift model for tasks like biomedical question answering (QA) and document classification using dense fine-tuning and soft prompting. We align tasks with varying output formats to our pre-training format by converting task labels into natural language sequences (Hu et al., 2021; Li and Liang, 2021). This method avoids structured formats and special tokens, ensuring semantic coherence and making full use of the natural language corpus MediSwift was trained on.
Following Luo et al. (2022), each downstream fine-tuning task is represented by a training dataset consisting of source-target pairs defined as: , where both and are sequences of tokens. For example, in question answering (e.g., PubMedQA), corresponds to the question and reference context description, and the categorical answer to the question; in biomedical document classification (e.g., Hallmarks of Cancers corpus), is the text passage and corresponds to the hallmarks of cancer.
@@ 2.2.2 Soft Prompting
Previous research on language models in the biomedical field has mainly focused on fine-tuning for domain-specific tasks (Bolton et al., 2021). Recently, there has been a move towards improving biomedical NLP task performance through prompt engineering (Luo et al., 2022; Nori et al., 2023b; Yagnik et al., 2024). Drawing inspiration from these advancements, our approach integrates prompt-based techniques into the fine-tuning phase of MediSwift. More precisely, we adopt the soft prompting methodology as described by Luo et al. (2022), aiming to refine our model’s capability in understanding and processing biomedical text. Similar to prefix-tuning (Li and Liang, 2021), to integrate soft prompts, we insert virtual tokens between the source and target sequences, thus modifying the loss function in Eq. 2. This adjustment accounts for the [source; prompt; target] sequence structure, impacting the model’s learning and inference. Let denote the prompt, consisting of a sequence of virtual tokens , where is the number of virtual tokens, and these tokens are represented by continuous embeddings. The modified Eq. 2, reflecting the inclusion of the prompt and its positioning, can be formalized as:
@@ 3.1 PubMed Papers and Abstracts
PMC consists of 4.91M full-text papers, and PubMed and PMC Abstracts comprise of 16.1 million papers (see Table 1). Moreover, we gathered the most recent PubMed entries, updated prior to 2023, directly from the official website, utilizing the official scripts for PubMed Abstracts222https://github.com/thoppe/The-Pile-PubMed?tab=readme-ov-file and PubMed Central333https://github.com/EleutherAI/pile-pubmedcentral. Similar to Luo et al. (2022) and Chen et al. (2023), we filter out empty items containing solely titles without accompanying abstracts.
Furthermore, prior works have show the significance of in-domain vocabulary for improving performance of specialized language models (Gu et al., 2021; Wu et al., 2023b; Mielke et al., 2021), a critical step that is often overlooked. Therefore, inspired by Luo et al. (2022) and (Bolton et al., 2021), instead of using the standard GPT-3 vocabulary, we learned the vocabulary directly from the biomedical corpus. Employing Moses (Koehn et al., 2007) tokenization followed by byte pair encoding (BPE), we segment the corpus into word pieces and learn the vocabulary; resulting in a size of 42,384. By exclusively pre-training with biomedical texts and using a specialized vocabulary, MediSwift improves the efficiency-accuracy frontier, as empirically shown through fine-tuning on biomedical tasks in Section 4.
@@ 3.2 MediSwift Pre-training
We pre-train and benchmarked MediSwift in-domain biomedical language models at 3 sizes: 302M, 510M and 1.21B. All MediSwift models are pre-trained from scratch using the Cerebras CS-2444https://docs.cerebras.net/en/2.1.1/wsc/how_to_guides/sparsity.html, taking advantage of its ability to accelerate training with unstructured sparsity. At the time of the study, the specialized kernels of Cerebras CS-2 were designed to facilitate training with static unstructured sparsity (refer to Appendix B for additional details regarding the hardware accelerator).
In the pre-training phase of the MediSwift models, a couple sparsity levels are explored, including 50%, and 75%, aside from their respective dense counterparts. The pre-training of MediSwift models are conducted on a single CS-2 for a total of 200,000 steps, with a batch size of 512 and a maximum sequence length of 1024 tokens, resulting in approximately 104.86B tokens processed in total (see Appendix A.1 for additional pre-training experimental setup details).
@@ 3.3 Sparse Pre-trained MediSwift
While there are several advanced sparse training techniques (Evci et al., 2020; Mocanu et al., 2018; Liu et al., 2021), for simplicity, in this work, we adopt static sparsity, namely random pruning, for the sparse pre-training of MediSwift models. This approach mandates a uniform distribution of sparsity levels throughout each layer, irrespective of the layer’s parameter count or its FLOPs. Specifically, the scope of our sparsification process is confined to all dense linear layers within the network, including both matrices within the multi-layer perceptron (MLP) module—namely, the intermediate layer and the MLP output projection, as well as the four weight matrices integral to the self-attention (Vaswani et al., 2017) mechanism: query, key, value, and attention output projection. Notably, we ensure that the embeddings, Layer Normalization (Ba et al., 2016) components and biases are kept dense.
@@ 3.4 Analysis on Pre-trained Models
In Figure 2, we illustrate the training loss curves for the MediSwift-XL model for both the dense and sparse configurations at . The 50% sparse model’s training loss closely follows that of the dense MediSwift-XL, showcasing minimal deviation throughout the training process. However, a noticeable divergence is observed at 75% sparsity, where the final training loss slightly lags behind that of the dense counterpart. Interestingly, when comparing the 75% sparse MediSwift-XL model with the dense MediSwift-Med model, the former, despite sharing the same number of non-embedding parameters, it achieves a lower training loss (refer to Table 2). This observation aligns with previous findings that larger, albeit sparser, models can surpass their smaller, densely parameterized equivalents in terms of performance (Thangarasa et al., 2023b; Liu et al., 2022a; Ramanujan et al., 2020; Golubeva et al., 2021). This superiority is further corroborated by the improved performance of the 75% sparse model on biomedical NLP tasks, when compared to the MediSwift-Med dense model (see Section 4), highlightling the benefits of training larger sparse models in comparison to smaller dense ones.
In Table 2, we provide an analysis of the computational efficiency achieved through sparse pre-training in our proposed MediSwift architectures, namely MediSwift-Med and MediSwift-XL. Here, we quantify the total FLOPs required for both forward and backward propagations during the pre-training phase of these models.
For the MediSwift-XL model, attention and vocabulary embeddings represent 13.3% and 6.8% of total FLOPs, respectively, hence highlighting the computational savings at the 1.2B parameter scale. Sparse pre-training with 75% sparsity reduces FLOPs by slightly over 2.5x compared to its dense counterpart. The smaller MediSwift-Med model has a higher FLOP percentage for attention and embeddings, hence achieving a 2x reduction at the same sparsity level. This indicates that FLOP savings increase with model size, demonstrating that larger models (Hoffmann et al., 2022; Kaplan et al., 2020; Hestness et al., 2017) can potentially benefit more from sparse pre-training. In addition, we emphasize that the total FLOPs required for fine-tuning these models account for a minor fraction of the overall pre-training FLOPs, reinforcing the efficiency of our approach in scaling to larger model sizes while conserving computational resources.

### v2

@@ Abstract
Vithursan Thangarasa† Mahmoud Salem Shreyas Saxena Kevin Leong Joel Hestness Sean Lie Cerebras Systems {vithu, sean}@cerebras.net
@@ 1 Introduction
Although domain-specific LLM pre-training offers significant benefits, the computational demands remain a significant challenge, necessitating innovative solutions for broader accessibility and efficiency. Techniques proposed to mitigate these burdens include sparse attention (Dao et al., 2022b; Jaszczur et al., 2021), quantized optimization (Tang et al., 2021), low-rank factorization (Lialin et al., 2023), and sequence-level curriculum learning (Li et al., 2022). Sparse attention enhances transformer efficiency but mainly benefits attention layers and may struggle with long-range dependencies. Quantized optimization reduces memory footprint and accelerates training with lower-precision arithmetic but introduces quantization noise and relies on hardware support. Low-rank factorization saves memory and computation by approximating weight matrices, risking information loss and requiring architectural changes. Sequence-level curriculum learning improves convergence and generalization by progressively increasing difficulty, primarily enhancing training efficiency. Among these approaches, weight sparsity emerges as a promising method, distinct from the others in its approach by setting a subset of model parameters to zero, thus reducing the FLOPs needed during training. Unlike sparse attention, which optimizes a specific component, or quantized optimization and low-rank factorization, which may introduce trade-offs in model performance, sparse pre-training directly reduces computational costs across the entire model without inherently compromising performance.
@@
However, the adoption of sparse training is limited by (1) the challenge of finding optimal sparsity patterns which retain the accuracy of the original dense model (Frankle and Carbin, 2018; Ma et al., 2022) and (2) its difficulties in accelerating on hardware optimized for dense computations (e.g., Nivida GPUs and Google TPUs) (Hooker, 2020). Additionally, sparse pre-training of LLMs typically leads to an accuracy loss from optimization challenges in sparse networks (Evci et al., 2019), but previous studies have shown that transitioning from sparse-to-dense training can effectively address these problems (Thangarasa et al., 2023a; Dao et al., 2022a). Our work aims to reduce training FLOPs by employing unstructured weight sparsity in domain-specific LLMs and recovering lost representational capacity by transitioning to dense weight matrices for downstream fine-tuning.
@@
Despite weight sparsity not being widely used in real-world applications, advancements in specialized software kernels (Gale et al., 2019; Elsen et al., 2019; Ashby et al., 2019; Tang et al., 2023) have facilitated inference acceleration with unstructured sparsity. Recent developments have shown that the benefits of unstructured weight sparsity can be fully harnessed on specialized hardware, such as the Cerebras CS-2 (Cerebras, 2023; Lie, 2022b), for LLM training. As sparse training techniques and hardware continue to co-evolve, we anticipate that the reductions in FLOPs will lead to realized sparse acceleration. The latest innovations in software and hardware (NeuralMagic, 2021; Gupta, 2024) are geared towards enabling the widespread adoption of unstructured weight sparsity, offering the potential to achieve higher compression ratios and practical speedups in terms of wall-clock time.
@@
Building on this momentum, we introduce MediSwift, a suite of biomedical language models (LMs) available in three sizes: Med (302M), Large (510M), and XL (1.21B). These models are based on GPT-3 and pre-trained sparsely from scratch on biomedical texts, aimed at reducing the computational costs required for training. We explore the impact of applying 50% and 75% weight sparsity during pre-training, which results in a 2-2.5x reduction in the overall training FLOPs needed. Figure 1 summarizes the performance, where our dense and sparse MediSwift models noticeably outperform other language models up to 7B parameters. We demonstrate MediSwift’s capabilities through fine-tuning on established benchmarks for biomedical natural language processing (NLP) tasks (e.g., PubMedQA (Jin et al., 2019) for question answering and HoC (Baker et al., 2016) for document classification) showing significant improvements in the balance between efficiency and accuracy.
Although previous research suggests that sparse pre-training may compromise model accuracy on downstream tasks, our approach incorporates dense fine-tuning with strategic soft prompting to effectively regain performance on specialized tasks. Specifically, MediSwift-XL (1.21B) sets a new state-of-the-art by reaching 76.8% accuracy, despite being 5.8x smaller than PMC-LlaMA. Moreover, MediSwift-XL models, pre-trained with 50% and 75% sparsity, surpass the performance of models of similar or greater sizes (e.g., MediSwift-XL at 75% sparsity outperforms BioMedLM while being almost 9x smaller). Our work not only highlights the potential for sparse pre-training to make LM training more economically viable but also sets a new benchmark for efficiency in domain-specific applications of LLMs. The key contributions are:
@@ 2.1 Autoregressive Language Modeling
Building upon our framework for dense pre-training, we introduce weight sparsity into the model, specifically to improve the computational efficiency. We achieve this by methodically reducing the number of active connections within each layer of the model by a predefined sparsity level , effectively rendering a portion of the network’s parameters inactive. This process yields a network with active parameters per layer, where denotes the original number of parameters in layer . The overall sparsity is quantified by , which represents the ratio of inactive parameters to the total parameter count of the initially dense model, calculated as . To apply sparsity effectively, we employ a binary mask to the model’s initial parameters , resulting in a sparse parameter set . Our approach to inducing sparsity involves random parameter pruning, a process where percentage of the model’s weights are randomly set to zero at initialization. This mask effectively segregates the parameters into active (1) and inactive (0) states, thereby establishing a sparsity-induced version of the language model that aims to minimize a slightly modified objective:
We leverage the GPT-3 (Brown et al., 2020a) architecture for the MediSwift biomedical language model, training it with the AdamW (Loshchilov and Hutter, 2017) optimizer on a curated biomedical dataset, following the objective shown in Eq. 1 for iterations to obtain parameters . This pre-trained model is then fine-tuned for specific supervised tasks in the biomedical domain. GPT-3 was selected for MediSwift due to its versatility in NLP tasks and suitability for specialized domains like biomedicine. Its autoregressive nature and capability to produce contextually relevant text align with our methodology of sparse pre-training, dense fine-tuning, and soft prompting to enhance efficiency and task performance. Additionally, GPT-3 was well-supported on the Cerebras CS-2 hardware, with optimized kernels for sparse training using unstructured weight sparsity, further enhancing our model’s computational efficiency.
Our approach is fundamentally model-agnostic and adaptable to various LLMs. The core components of our methodology (i.e., sparse pre-training to reduce computational burden, followed by dense fine-tuning and soft prompting to regain or enhance task performance) can be broadly applied across different LLMs, including newer iterations of GPT.
@@ 2.2 Dense Fine-tuning and Soft Prompting
In this section, we detail the adaptation of our pre-trained MediSwift model for tasks like biomedical question answering (QA) and document classification using dense fine-tuning and soft prompting. We align tasks with varying output formats to our pre-training format by converting task labels into natural language sequences (Li and Liang, 2021; Hu et al., 2021). This method avoids structured formats and special tokens, ensuring semantic coherence and making full use of the natural language corpus MediSwift was trained on.
Following Luo et al. (2022), each downstream fine-tuning task is represented by a training set consisting of source-target pairs defined as: , where both and are sequences of tokens. For example, in question answering (e.g., PubMedQA), corresponds to the question and reference context description, and the categorical answer to the question; in biomedical document classification (e.g., Hallmarks of Cancers corpus), is the text passage and corresponds to the hallmarks of cancer.
@@ 2.2.2 Soft Prompting
Previous research on language models in the biomedical field has mainly focused on fine-tuning for domain-specific tasks (Bolton et al., 2021). Recently, there has been a move towards improving biomedical NLP task performance through prompt engineering (Luo et al., 2022; Nori et al., 2023b; Yagnik et al., 2024). Drawing inspiration from these advancements, our approach integrates prompt-based techniques into the fine-tuning phase of MediSwift. More precisely, we adopt the soft prompting methodology as described by Luo et al. (2022), aiming to refine our model’s capability in understanding and processing biomedical text. Similar to existing work on prompt tuning (Brown et al., 2020b; Liu et al., 2021b; Lester et al., 2021), to integrate soft prompts, we insert virtual tokens between the source and target sequences, thus modifying the loss function in Eq. 2. This adjustment accounts for the [source; prompt; target] sequence structure, impacting the model’s learning and inference. Let denote the prompt, consisting of a sequence of virtual tokens , where is the number of virtual tokens, and these tokens are represented by continuous embeddings. The modified Eq. 2, reflecting the inclusion of the prompt and its positioning, can be formalized as:
@@ 3.1 PubMed Papers and Abstracts
PMC consists of 4.91M full-text papers, and PubMed and PMC Abstracts comprise of 16.1 million papers (see Table 1). Moreover, we gathered the most recent PubMed entries, updated prior to 2023, directly from the official website, utilizing the official scripts for PubMed Abstracts222https://github.com/thoppe/The-Pile-PubMed?tab=readme-ov-file and PubMed Central333https://github.com/EleutherAI/pile-pubmedcentral. Similar to Luo et al. (2022) and Chen et al. (2023), we filter out empty items containing solely titles without accompanying abstracts. Furthermore, prior works have show the significance of in-domain vocabulary for improving performance of specialized LMs (Gu et al., 2021; Wu et al., 2023b; Mielke et al., 2021), a critical step that is often overlooked. Therefore, inspired by Luo et al. (2022) and (Bolton et al., 2021), instead of using the standard GPT-3 vocabulary, we learned the vocabulary directly from the biomedical corpus. Employing Moses (Koehn et al., 2007) tokenization followed by byte pair encoding (BPE), we segment the corpus into word pieces and learn the vocabulary; resulting in a size of 42,384. By exclusively pre-training with biomedical texts and using a specialized vocabulary, MediSwift improves the efficiency-accuracy frontier, as empirically shown on biomedical tasks in Section 4.
@@ 3.2 MediSwift Pre-training
We pre-train and benchmarked MediSwift in-domain biomedical language models at 3 sizes: 302M, 510M and 1.21B. All MediSwift models are pre-trained from scratch using the Cerebras CS-2444https://docs.cerebras.net/en/2.1.1/wsc/how_to_guides/sparsity.html, taking advantage of its ability to accelerate training with unstructured sparsity. At the time of the study, the specialized kernels of Cerebras CS-2 were designed to facilitate training with static unstructured sparsity (refer to Appendix B for additional details). In the pre-training phase of the MediSwift models, 50%, and 75% sparsity levels are explored, aside from their respective dense counterparts. The pre-training of MediSwift models are conducted on a single CS-2 for a total of 200,000 steps, with a batch size of 512 and a maximum sequence length of 1024 tokens, resulting in approximately 104.86B tokens processed in total (see Appendix A.1 for additional pre-training experimental setup details).
@@ 3.3 Sparse Pre-trained MediSwift
While there are several advanced sparse training techniques (Evci et al., 2020; Mocanu et al., 2018; Liu et al., 2021a), for simplicity, in this work, we adopt static sparsity, namely random pruning, for the sparse pre-training of MediSwift models. This approach mandates a uniform distribution of sparsity levels throughout each layer, irrespective of the layer’s parameter count or its FLOPs. Specifically, the scope of our sparsification process is confined to all dense linear layers within the network, including both matrices within the multi-layer perceptron (MLP) module—namely, the intermediate layer and the MLP output projection, as well as the four weight matrices integral to the self-attention (Vaswani et al., 2017) mechanism: query, key, value, and attention output projection. Notably, we ensure that the embeddings, Layer Normalization (Ba et al., 2016) components and biases are kept dense.
@@ 3.4 Analysis on Pre-trained Models
In Figure 2, we illustrate the training loss curves for the MediSwift-XL model for both the dense and sparse configurations at . The 50% sparse model’s training loss closely follows that of the dense MediSwift-XL, showcasing minimal deviation throughout the training process. However, a noticeable divergence is observed at 75% sparsity, where the final training loss slightly lags behind that of the dense counterpart. Interestingly, when comparing the 75% sparse MediSwift-XL model with the dense MediSwift-Med model, the former, despite sharing the same number of non-embedding parameters, it achieves a lower training loss (refer to Table 2). This observation aligns with previous findings that larger, albeit sparser, models can surpass their smaller, densely parameterized equivalents in terms of performance (Thangarasa et al., 2023b; Liu et al., 2022a; Ramanujan et al., 2020; Golubeva et al., 2021). This superiority is further supported by the improved accuracy of the 75% sparse model on biomedical NLP tasks, when compared to the MediSwift-Med dense model (see Section 4), highlightling the benefits of training larger sparse models in comparison to smaller dense ones.
In Table 2, we provide an analysis of the computational efficiency achieved through sparse pre-training in our proposed MediSwift architectures, namely MediSwift-Med and MediSwift-XL. Here, we quantify the total FLOPs required for both forward and backward propagations during the pre-training phase of these models. For the MediSwift-XL model, attention and vocabulary embeddings represent 13.3% and 6.8% of total FLOPs, respectively, hence highlighting the computational savings at the 1.21B parameter scale. Sparse pre-training with 75% sparsity reduces FLOPs by slightly over 2.5x compared to its dense counterpart. The smaller MediSwift-Med model has a higher FLOP percentage for attention and embeddings, hence achieving a 2x reduction at the same sparsity level. This indicates that FLOP savings increase with model size, demonstrating that larger models (Hoffmann et al., 2022; Kaplan et al., 2020; Hestness et al., 2017) can potentially benefit more from sparse pre-training. In addition, we emphasize that the total FLOPs required for fine-tuning these models account for a minor fraction of the overall pre-training FLOPs, reinforcing the efficiency of our approach in scaling to larger model sizes while conserving computational resources.